APPLICATIONS OF EFFECTIVE ROTATIONAL HAMILTONIANS FOR TWO-ROTOR MOLECULES

\underline{Peter Groner}, Department of Chemistry, University of Missouri - Kansas City, Kansas City, MO 64110.

Effective rotational Hamiltonians for molecules with one periodic internal motion have been described a few years ago ^{a}. The same paper also outlined the extension of the model to molecules with more than one periodic motions. Such a Hamiltonian was recently implemented in a computer program to analyze spectra of tworotor molecules such as CH_{3}OCH_{3} and CH_{3}COCH_{3}. Because the analysis was not completely satisfactory, the theoretical treatment was modified producing a different effective rotational Hamiltonian for such molecules. It was subsequently applied, with better success, to the analysis of the same spectra. For the ground state of CH_{3}OCH_{3}, 232 rotational transition frequencies of all symmetry species taken from the literature could be fit to an effective rotational Hamiltonian containing 16 parameters, among them 4 internal rotation or tunneling parameters, to yield a standard deviation of 60 kHz; however, about 20 assigned transition frequencies had to be excluded from the fit. Detailed results for the ground and vibrational excited states of CH_{3}OCH_{3}, for the ground state of CH_{3}COCH_{3}, and for other examples, will be given in the presentation.

aP. Groner, J. Mol. Spectrosc. 156, 164-189 (1992).